# Local-global compatibility and the exceptional zero conjecture for $\operatorname{GL}(3)$

### v1

@@ 4.2. $p$-adic measures attached to $p$-arithmetic cohomology classes
$\Big{(}u_{0}\cdot\varphi_{\mathcal{A},\mathcal{B},f}\Big{)}(g)=\left\{\begin{array}[]{cc}f(y)&\text{ if }g\in\overline{B}(\mathbb{Q}_{p})u_{0}^{-1}\left(\begin{smallmatrix}1&x&y\\ &1&\\ &z&1\end{smallmatrix}\right)\text{ with }x\in\mathcal{A},\ y\in\mathcal{B},\ z\in\mathbb{Z}_{p}\\ 0&\text{ otherwise, }\end{array}\right$.
@@ 8.1. Families
$A^{r\operatorname{-an}}_{\Omega}=\{f\colon\operatorname{Iw}\to\mathcal{O}(\Omega):f\text{ is }r\text{-analytic},f(b\cdot-)=\kappa_{\Omega}(b)\cdot f(-)\;\forall b\in\overline{B}(\mathbb{Q}_{p})\cap\operatorname{Iw}\}$
and let $D^{r\operatorname{-an}}_{\Omega}$ denote the continuous dual of $A^{r\operatorname{-an}}_{\Omega}$. Let $T^{+}$ (resp. $T^{-}$) be the submonoid of $T(\mathbb{Q}_{p})$ of elements which have non-negative valuation when paired with positive (resp. negative) roots. The space $A^{r\operatorname{-an}}_{\Omega}$ carries actions of $\operatorname{Iw}$ and $T^{-}$, and $D^{r\operatorname{-an}}_{\Omega}$ carries actions of $\operatorname{Iw}$ and $T^{+}$, defined for our normalisations in [BW21, §3.4]. Note the actions of $T^{\pm}$ and $\operatorname{Iw}$ agree on the intersections.
@@ 8.3. The Galois Benois–Colmez–Greenberg–Stevens formula
In the introduction, this assumption is built into the ‘Steinberg-at-$p$’ assumption on $\rho_{\pi}$.
@@ 9.2.2. Characteristic $p$ results
Let $k=\mathcal{O}/\varpi$, and let $\mathfrak{n}$ be a maximal ideal in the (prime-to-$S$) Hecke algebra $\widetilde{\mathbf{T}}^{S}\left(\operatorname{H}^{*}(X_{\widetilde{G},K},k)\right)$ (resp. $\mathbf{T}^{S}\left(H^{*}(X_{\operatorname{GL}_{n},K},k\right)$) for some sufficiently small compact open subgroup $K\subset\widetilde{G}(\mathbb{A}_{f})$ (resp. $K\subset\operatorname{GL}_{n}(\mathbb{A}_{f})$). There exists a continuous semisimple Galois representation
@@
For the standard representation for general symplectic groups, the existence is explained in the proof of [Sch15, Corollary 5.2.6] (and relies on Arthur’s endscopic classification for symplectic groups [Art13]). The existence for general linear groups again follows from [Sch15] (which, in fact, uses the existence for general symplectic groups as input). ∎
@@ 9.3. Vanishing results and boundary cohomology
We now collect together key cohomological results; firstly, a vanishing result of Yang and Zhu in characteristic $p$, and secondly, an analysis of boundary cohomology that will allow the transfer of classes from $\widetilde{G}$ to $M$ via a degree-shifting argument.
@@ 9.3.1. Vanishing results
Recall that $L/\mathbb{Q}_{p}$ is a finite extension with ring of integers $\mathcal{O}$ and unformiser $\varpi$. Let $\mathfrak{m}\subset\mathbf{T}^{S}$ be a maximal ideal with residue field $k=\mathcal{O}/\varpi$, and let $\Theta\colon\mathbf{T}^{S}\twoheadrightarrow k$ denote the corresponding $\mathcal{O}$-algebra morphism with kernel $\mathfrak{m}$.
Assumption 9.3.1.
Let $\widetilde{\mathfrak{m}}=S^{*}(\mathfrak{m})$ denote the pullback of $\mathfrak{m}$ under the unnormalised Satake morphism. We set
@@ 9.5. Transfer map on Hecke algebras
Let $p>2n$. Let $m\geqslant 1$ and $c\geqslant b\geqslant 0$ with $c\geqslant 1$, and let $\widetilde{K}$ and $K$ be as in §9.3.3 with the additional condition that $\widetilde{K}_{p}=\widetilde{\operatorname{Iw}}(b,c)$ and $K_{p}=\operatorname{Iw}(b,c)$. Set $d=n(n+1)/2$ for brevity. Let $\widetilde{\lambda}$ be a dominant weight for $\widetilde{G}$. We assume that $\overline{\rho}_{\widetilde{\mathfrak{m}}}$ is decomposed generic. From the excision exact sequence, we obtain a Hecke-equivariant map
@@
Combining (9.5.1), (9.5.5) and Lemma 9.5.6, we obtain a subquotient
$\mathrm{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda}}/\varpi^{m})^{\operatorname{ord}}_{\widetilde{\mathfrak{m}}}$$\operatorname{H}^{d-l(x)}(X_{M,K},V_{\lambda_{x}}/\varpi^{m})^{\operatorname{ord}}_{\mathfrak{m}}.$}} (9.5.7)
This is $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}$-equivariant, where the action on the target is via $\mathcal{S}_{w}\colon\widetilde{\mathbf{T}}^{S,\operatorname{ord}}\to\mathbf{T}^{S,\operatorname{ord}}$; recall $\mathcal{S}_{w}$ is the ring homomorphism coinciding with $\mathcal{S}$ away from $p$, and which at $p$ is induced from the map
@@
With notation as above, the ordinary part $\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}$ carries a Hecke-equivariant descending filtration
$\operatorname{Fil}^{i}\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}$
@@
  • $\operatorname{Fil}^{0}\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}=\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}$ and $\operatorname{Fil}^{d+1}\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}=0$;
@@ 9.6.1. Characteristic zero local-global compatibility
Recall that $p>2n$ and $\mathfrak{m}$ is assumed to be decomposed generic and non-Eisenstein, and we assume that $\widetilde{\mathfrak{m}}$ is decomposed generic. Let $c\geqslant b\geqslant 0$ with $c\geqslant 1$, and let $\widetilde{K}$ and $K$ be as in §9.3.3 with the additional condition that $\widetilde{K}_{p}=\widetilde{\operatorname{Iw}}(b,c)$ and $K_{p}=\operatorname{Iw}(b,c)$.
@@
    equals the image of $\widetilde{H}_{\ell}(X)$ in $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}}[1/p])_{\mathfrak{n}}[X]$. Moreover, the residual representation $\overline{\rho}_{\mathfrak{n}}$ modulo $p$ is isomorphic to $\overline{\rho}_{\widetilde{\mathfrak{m}}}$.
@@
    where $\psi_{\widetilde{\lambda},i}\colon G_{\mathbb{Q}_{p}}\to\overline{\mathbb{Q}}_{p}^{\times}$ is the character as in Definition 9.1.2 associated with the algebra homomorphism $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}\to\overline{\mathbb{Q}}_{p}$ with kernel $\mathfrak{n}$.
@@
Recall that we are assuming $n\geqslant 2$. The torsion subgroup $\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}[\varpi]$ is a quotient of $\operatorname{H}^{d-1}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda}}/\varpi)_{\widetilde{\mathfrak{m}}}$, which is zero by Theorem 9.3.2. Fix an isomorphism $\mathbb{C}\cong\overline{\mathbb{Q}}_{p}$. Via this isomorphism, we can consider the singular cohomology
@@
Now let $\mathfrak{n}$ be a maximal ideal as in the statement. Base-changing to $\overline{\mathbb{Q}}_{p}$, we obtain a maximal ideal $\mathfrak{b}\subset\widetilde{\mathbf{T}}^{S}_{\overline{\mathbb{Q}}_{p}}$ lying above $\mathfrak{n}$ and $\widetilde{\mathfrak{m}}$. Theorem 9.3.2 implies $\operatorname{H}^{d+1}_{c}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}=0,$ so the same is true after inverting $p$ and localising at $\mathfrak{b}$. Thus (9.6.2) induces a Hecke-equivariant isomorphism
$\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})_{\mathfrak{b}}\cong\operatorname{H}^{d}_{\operatorname{cusp}}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})_{\mathfrak{b}}\oplus\operatorname{H}^{d}(\partial\overline{X}_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})_{\mathfrak{b}}$ (9.6.3)
(and similarly after taking ordinary parts). We see that if $\mathfrak{b}$ is a maximal ideal in $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})^{\operatorname{ord}}),$ then it contributes to at least one of cuspidal or boundary cohomology. Suppose first it contributes to the cuspidal part; then by the usual description of cuspidal cohomology, $\mathfrak{b}$ is the prime-to-$S$ Hecke eigensystem associated with a regular algebraic cuspidal automorphic representation $\sigma$ of $\widetilde{G}(\mathbb{A})$ of weight $w_{0}^{\widetilde{G}}\widetilde{\lambda}$ which is ordinary at $p$, and we may take $\rho_{\mathfrak{n}}=\rho_{\sigma}$ via Theorem 9.2.4.
@@
for some $w\in{{}^{M}W}$ and $\mathfrak{a}$ as above (where the second summand is zero if no such $w$ exists). Thus $\mathfrak{b}$ arises from a regular algebraic cuspidal automorphic representation $\sigma$ of $\widetilde{G}(\mathbb{A})$ (if it contributes to the first summand) or $M(\mathbb{A})$ (the second). As the ordinary part of $\sigma$ is at most 1-dimensional, we see that $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}$ acts semisimply on $\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})^{\operatorname{ord}}_{\mathfrak{b}}$, completing the proof of (1).
@@
By construction this satisfies the properties desired in (2). It remains to deduce (3) in this case. As explained above, the Galois representation $\rho_{\sigma}$ satisfies local-global compatibility at $\ell=p$ by Theorem 9.2.6. Since $w\in{{}^{M}W}$, we may re-order the filtration on $\rho_{\mathfrak{n}}|_{G_{\mathbb{Q}_{p}}}$ so that all the desired properties are satisfied. More precisely, the parameters
@@
for $\rho_{\mathfrak{n}}$ (excluding the trivial parameter) are in order
@@
in order. Here $\chi_{\lambda_{x},i}$ are the characters associated with $\sigma$ as in Definition 9.1.2. ∎
We note the following important consequence of this proposition:
@@
Consider notation and assumptions as in Proposition 9.6.1. There exists a continuous $(2n+1)$-dimensional determinant
@@
such that for any $\mathfrak{n}$ as in Proposition 9.6.1(2), we have $\Theta_{\mathfrak{n}}\circ\widetilde{D}=\operatorname{det}(\rho_{\mathfrak{n}})$ as maps $G_{\mathbb{Q}}\to\overline{\mathbb{Q}}_{p}$. Here $\Theta_{\mathfrak{n}}\colon\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}})\to\overline{\mathbb{Q}}_{p}$ is reduction modulo $\mathfrak{n}$.
@@
This follows from [Che14, Example 2.32] applied to the embedding
$\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}})\hookrightarrow\prod_{\mathfrak{n}}\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}}[1/p])_{\mathfrak{n}}.\qed$

### v2

@@ 4.2. $p$-adic measures attached to $p$-arithmetic cohomology classes
[…]
@@ 8.1. Families
$A^{r\operatorname{-an}}_{\Omega}=\{f\colon\operatorname{Iw}\to\mathcal{O}(\Omega):f\text{ is }r\text{-analytic},f(b\cdot-)=\kappa_{\Omega}(b)\cdot f(-)\;\forall b\in\overline{B}(\mathbb{Q}_{p})\cap\operatorname{Iw}\},$
where a function is $r$-analytic if it is analytic on each open disc of radius $p^{-r}$ in $\operatorname{Iw}$ (see e.g. [BW21, §3.2.2]), and let $D^{r\operatorname{-an}}_{\Omega}$ denote the continuous $\mathcal{O}(\Omega)$-dual of $A^{r\operatorname{-an}}_{\Omega}$. Let $T^{+}$ (resp. $T^{-}$) be the submonoid of $T(\mathbb{Q}_{p})$ of elements which have non-negative valuation when paired with positive (resp. negative) roots. The space $A^{r\operatorname{-an}}_{\Omega}$ carries left-actions of $\operatorname{Iw}$ (by right translation) and $T^{-}$; the latter is induced by the map $N(\mathbb{Z}_{p})\to N(\mathbb{Z}_{p})$, $n\mapsto[t]t^{-1}nt$, where $[t]$ is the image of $t$ under the natural projection $T(\mathbb{Q}_{p})\twoheadrightarrow T(\mathbb{Z}_{p})$ (cf. [BW21, §3.4]). Dualising, $D^{r\operatorname{-an}}_{\Omega}$ carries left actions of $\operatorname{Iw}$ and $T^{+}$. Note the two actions agree for any $t\in T^{\pm}\cap\operatorname{Iw}$.
@@ 9.1.6. Borel–Serre compactifications
9.1.7. Running set-up
We finally fix our running set-up. Recall that $L/\mathbb{Q}_{p}$ is a finite extension with ring of integers $\mathcal{O}$ and unformiser $\varpi$. Let $\mathfrak{m}\subset\mathbf{T}^{S}$ be a maximal ideal with residue field $k=\mathcal{O}/\varpi$, and let $\Theta\colon\mathbf{T}^{S}\twoheadrightarrow k$ denote the corresponding $\mathcal{O}$-algebra morphism with kernel $\mathfrak{m}$.
Assumption 9.1.6.
Let $\widetilde{\mathfrak{m}}=\mathcal{S}^{*}(\mathfrak{m})\subset\widetilde{\mathbf{T}}^{S}$ be the pullback of $\mathfrak{m}$ under the unnormalised Satake morphism. We set
@@ 9.2.2. Characteristic $p$ results
Let $m\geqslant 1$ be an integer, and let $\mathfrak{n}$ be a maximal ideal with finite residue field in the (prime-to-$S$) Hecke algebra $\widetilde{\mathbf{T}}^{S}\left(\operatorname{H}^{*}(X_{\widetilde{G},K},\mathcal{O}/\varpi^{m})\right)$ (resp. $\mathbf{T}^{S}\left(\operatorname{H}^{*}(X_{\operatorname{GL}_{n},K},\mathcal{O}/\varpi^{m})\right)$) for some sufficiently small compact open subgroup $K\subset\widetilde{G}(\mathbb{A}_{f})$ (resp. $K\subset\operatorname{GL}_{n}(\mathbb{A}_{f})$). Then there exists a continuous semisimple Galois representation
@@
For the standard representation for general symplectic groups, the existence is explained in the proof of [Sch15, Corollary 5.2.6] (and relies on Arthur’s endscopic classification for symplectic groups [Art13]). The existence for general linear groups again follows from [Sch15] (which, in fact, uses the existence for general symplectic groups as input). In both cases, Scholze constructs a continuous determinant with the desired properties, valued in a quotient of the Hecke algebra; further quotienting by $\mathfrak{n}$, one gets determinants valued in $\overline{\mathbb{F}}_{p}$, and then concludes by [Che14, Thm. 2.12]. ∎
@@ 9.3. Vanishing results and boundary cohomology
We now collect together key cohomological results; firstly, a vanishing result of Yang and Zhu in characteristic $p$, and secondly, an analysis of boundary cohomology that will allow the transfer of classes from $\widetilde{G}$ to $M$ via a degree-shifting argument. Throughout, $\mathfrak{m}$ and $\widetilde{\mathfrak{m}}$ are the maximal ideals from §9.1.7; in particular, $\mathfrak{m}$ satisfies Assumption 9.1.6.
@@ 9.5. Transfer map on Hecke algebras
Let $p>2n$. Let $m\geqslant 1$ and $c\geqslant b\geqslant 0$ with $c\geqslant 1$, and let $\widetilde{K}$ and $K$ be as in §9.3.3 with the additional condition that $\widetilde{K}_{p}=\widetilde{\operatorname{Iw}}(b,c)$ and $K_{p}=\operatorname{Iw}(b,c)$. Set $d=n(n+1)/2$ for brevity. Let $\widetilde{\lambda}$ be a dominant weight for $\widetilde{G}$. Recall $\mathfrak{m}$ and $\widetilde{\mathfrak{m}}$ from §9.1.7; we assume that $\overline{\rho}_{\widetilde{\mathfrak{m}}}$ (also defined in §9.1.7) is decomposed generic. From the excision exact sequence, we obtain a Hecke-equivariant map
@@
Recall that $\tau_{w}^{-1}$ does not affect the underlying space; it means only that the $T(\mathbb{Q}_{p})$-action is twisted through the map $t\mapsto wtw^{-1}$. In particular, combining (9.5.1), (9.5.5) and Lemma 9.5.6, we obtain a subquotient
$\mathrm{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda}}/\varpi^{m})^{\operatorname{ord}}_{\widetilde{\mathfrak{m}}}$$\operatorname{H}^{d-l(x)}(X_{M,K},V_{\lambda_{x}}/\varpi^{m})^{\operatorname{ord}}_{\mathfrak{m}},$}} (9.5.7)
which is $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}$-equivariant if $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}$ acts on the source in the usual way, and on the target via $\mathcal{S}_{w}\colon\widetilde{\mathbf{T}}^{S,\operatorname{ord}}\to\mathbf{T}^{S,\operatorname{ord}}$. Here recall $\mathcal{S}_{w}$ is the ring homomorphism coinciding with $\mathcal{S}$ away from $p$, and which at $p$ is induced from the map
@@
With notation as above, the ordinary part $\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda}})^{\operatorname{ord}}$ carries a Hecke-equivariant descending filtration
$\operatorname{Fil}^{i}\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda}})^{\operatorname{ord}}$
@@
  • $\operatorname{Fil}^{0}\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda}})^{\operatorname{ord}}=\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda}})^{\operatorname{ord}}$ and $\operatorname{Fil}^{d+1}\operatorname{H}^{d}(X_{\widetilde{K}}^{P},V_{\widetilde{\lambda}})^{\operatorname{ord}}=0$;
@@ 9.6.1. Characteristic zero local-global compatibility
Recall that $p>2n$ and let $\mathfrak{m}\subset\mathbf{T}^{S}$ and $\widetilde{\mathfrak{m}}\subset\widetilde{\mathbf{T}}^{S}$ be as in §9.1.7. Recall $\mathfrak{m}$ is assumed to be decomposed generic and non-Eisenstein, and we assume that $\widetilde{\mathfrak{m}}$ is decomposed generic. Let $c\geqslant b\geqslant 0$ with $c\geqslant 1$, and let $\widetilde{K}$ and $K$ be as in §9.3.3 with the additional condition that $\widetilde{K}_{p}=\widetilde{\operatorname{Iw}}(b,c)$ and $K_{p}=\operatorname{Iw}(b,c)$.
@@
    equals the image of $\widetilde{H}_{\ell}(X)$ in $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}}[1/p])_{\mathfrak{n}}[X]$. Moreover, the residual representation $\overline{\rho}_{\mathfrak{n}}$ is isomorphic to $\overline{\rho}_{\widetilde{\mathfrak{m}}}$.
@@
    where $\psi_{\widetilde{\lambda},i}\colon G_{\mathbb{Q}_{p}}\to\overline{\mathbb{Q}}_{p}^{\times}$ is the character as in Definition 9.1.2, associated with the algebra homomorphism $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}_{L}\to\overline{\mathbb{Q}}_{p}$ with kernel $\mathfrak{n}$.
@@
Recall that we are assuming $n\geqslant 2$. The torsion subgroup $\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}[\varpi]$ is a quotient of $\operatorname{H}^{d-1}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda}}/\varpi)_{\widetilde{\mathfrak{m}}}$, which is zero by Theorem 9.3.1; thus the cohomology is $\varpi$-torsion-free.
Fix an isomorphism $\mathbb{C}\cong\overline{\mathbb{Q}}_{p}$. Via this isomorphism, we can consider the singular cohomology
@@
Now take ordinary parts and let $\mathfrak{n}$ be a maximal ideal as in the statement. Base-changing to $\overline{\mathbb{Q}}_{p}$, we obtain a maximal ideal $\mathfrak{b}\subset\widetilde{\mathbf{T}}^{S}_{\overline{\mathbb{Q}}_{p}}$ lying above $\mathfrak{n}$ and $\widetilde{\mathfrak{m}}$. Theorem 9.3.1 implies $\operatorname{H}^{d+1}_{c}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}_{\widetilde{\mathfrak{m}}}=0,$ so the same is true after inverting $p$ and localising at $\mathfrak{b}$; so (9.6.2) yields a short exact sequence $0\to(\mathrm{H}^{d}_{\operatorname{cusp}})^{\operatorname{ord}}_{\mathfrak{b}}\to(\mathrm{H}^{d})^{\operatorname{ord}}_{\mathfrak{b}}\to(\mathrm{H}^{d}_{\partial})^{\operatorname{ord}}_{\mathfrak{b}}\to 0$. This splits by [FS98, (7)], giving a Hecke-equivariant isomorphism
$\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})^{\operatorname{ord}}_{\mathfrak{b}}\cong\operatorname{H}^{d}_{\operatorname{cusp}}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})^{\operatorname{ord}}_{\mathfrak{b}}\oplus\operatorname{H}^{d}(\partial\overline{X}_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})^{\operatorname{ord}}_{\mathfrak{b}}.$ (9.6.3)
If $\mathfrak{b}$ contributes to $\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})^{\operatorname{ord}}$, it thus contributes to at least one of cuspidal or boundary cohomology. Suppose first it contributes to the cuspidal part; then by the usual description of cuspidal cohomology, $\mathfrak{b}$ is the prime-to-$S$ Hecke eigensystem associated with a regular algebraic cuspidal automorphic representation $\sigma$ of $\widetilde{G}(\mathbb{A})$ of weight $w_{0}^{\widetilde{G}}\widetilde{\lambda}$ which is ordinary at $p$, and we may take $\rho_{\mathfrak{n}}=\rho_{\sigma}$ via Theorem 9.2.4.
@@
for some $w\in{{}^{M}W}$ and $\mathfrak{a}$ as above (where the second summand is zero if no such $w$ exists). Thus $\mathfrak{b}$ arises from a regular algebraic cuspidal automorphic representation $\sigma$ of $\widetilde{G}(\mathbb{A})$ (if it contributes to the first summand) or $M(\mathbb{A})$ (the second). As the ordinary part of $\sigma$ is at most 1-dimensional, we see that $\widetilde{\mathbf{T}}^{S,\operatorname{ord}}$ acts semisimply on $\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\overline{\mathbb{Q}}_{p}})^{\operatorname{ord}}_{\mathfrak{b}}$, hence on $\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}}[1/p]$, completing the proof of (1).
@@
By construction this satisfies the properties desired in (2). It remains to deduce (3) in this case. As explained above, the Galois representation $\rho_{\sigma}$ satisfies local-global compatibility at $\ell=p$ by Theorem 9.2.6. Since $w\in{{}^{M}W}$, we may re-order the filtration on $\rho_{\mathfrak{n}}|_{G_{\mathbb{Q}_{p}}}$ so that all the desired properties are satisfied. More precisely the ordered set of parameters
@@
for $\rho_{\mathfrak{n}}$ (excluding the trivial parameter) is equal to the ordered set
@@
as ordered sets. Here $\chi_{\lambda_{x},i}$ are the characters associated with $\sigma$ as in Definition 9.1.2. ∎
We note two important consequences of this proposition. After inverting $p$, we obtain a Galois representation attached to $\widetilde{\mathfrak{m}}$ valued in the Hecke algebra satisfying local-global compatibility at $\ell\not\in S$ and at $p$. Integrally, whilst we don’t have a full representation, we can construct a Galois determinant.
@@
Consider notation and assumptions as in Proposition 9.6.1.
  • (1)
    After inverting $p$, there exists a continuous semisimple Galois representation
    $\sigma_{\widetilde{\mathfrak{m}}}:G_{\mathbb{Q}}\to\operatorname{GL}_{2n+1}\Big(\widetilde{\mathbf{T}}^{S,\operatorname{ord}}\Big(\mathrm{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}_{\widetilde{\mathfrak{m}}}\big[\tfrac{1}{p}\big]\Big)\Big)$
    such that
    • (a)
[…]
    • (b)
      one has
      $\sigma_{\widetilde{\mathfrak{m}}}|_{G_{\mathbb{Q}_{p}}}\sim\left(\begin{array}[]{cccc}\psi_{\widetilde{\lambda},1}&*&*&*\\ &\psi_{\widetilde{\lambda},2}&*&*\\ &&\ddots&\vdots\\ &&&\psi_{\widetilde{\lambda},2n+1}\end{array}\right),$
      where $\psi_{\widetilde{\lambda},i}\colon G_{\mathbb{Q}_{p}}\to\widetilde{\mathbf{T}}^{S,\operatorname{ord}}\Big(\mathrm{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})^{\operatorname{ord}}_{\widetilde{\mathfrak{m}}}\big[\tfrac{1}{p}\big]\Big)^{\times}$ is as in Definition 9.1.2.
  • (2)
    Integrally, there exists a continuous $(2n+1)$-dimensional determinant
@@
[…]
@@
For (1), we take $\sigma=\oplus_{\mathfrak{n}}\rho_{\mathfrak{n}}$, where $\mathfrak{n}$ ranges over the maximal ideals as Proposition 9.6.1.
For (2), by Proposition 9.6.1(1/2), we have an embedding
$\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}})\hookrightarrow\prod_{\mathfrak{n}}\widetilde{\mathbf{T}}^{S,\operatorname{ord}}(\operatorname{H}^{d}(X_{\widetilde{G},\widetilde{K}},V_{\widetilde{\lambda},\mathcal{O}})_{\widetilde{\mathfrak{m}}}^{\operatorname{ord}}[1/p])_{\mathfrak{n}}.$
The result then follows from [Che14, Example 2.32]. ∎